# CopyBench: Measuring Literal and Non-Literal Reproduction of Copyright-Protected Text in Language Model Generation

### v1

@@ Abstract
Tong Chen1Akari Asai1*Niloofar Mireshghallah1* Sewon Min1James Grimmelmann2Yejin Choi1,3 Hannaneh Hajishirzi1,3Luke Zettlemoyer1Pang Wei Koh1,3 1University of Washington   2Cornell University   3Allen Institute for AI
@@ 1 Introduction
To bridge the gap, we introduce CopyBench, a new benchmark and automatic evaluation protocols designed to assess reproduction of copyright-protected text by LMs (Figure 1). We evaluate two categories of copying: literal and non-literal copying. Literal copying assesses the extent to which a model can reproduce copyright-protected content exactly as it appears in the source material. In contrast, non-literal copying evaluates whether a model generates outputs that, despite differing in surface form (e.g., through paraphrasing), exhibit a high degree of overlap in content. To the best of our knowledge, our work is the first that evaluates non-literal reproduction of copyrighted work in language model generation. In order to study the trade-offs between the unintended copying and the desired utilities of LMs, we also quantify two aspects of desired utilities: fact recall, i.e., answering questions about book content, and fluency. Our benchmark therefore allows for the evaluation of levels of copyright work reproduction, overall model utility, and any associated trade-offs. We curate an dataset using a list of popular copyright-protected fictions, sourced from the famous CliffsNotes study guides, [^1] https://www.cliffsnotes.com/ which provide human-written plot summary for each book.
@@
Additionally, our datasets is designed to benchmark methods for potentially reducing copying behavior, broadly categorized into training methods (instruction tuning and chat alignment) and inference methods (e.g., MemFree decoding; Ippolito et al. 2023). We find that different instruction-tuning methods have varying levels of effectiveness. Specifically, Llama2-Chat Touvron et al. (2023) and Llama3-Instruct AI@Meta (2024) significantly reduce copying behavior, though the mechanism remains unclear due to the use of closed-source data. In contrast, the open-source model Tulu2 Ivison et al. (2023), which is based on Llama2 and further trained with fully open-sourced instruction tuning and preference data, shows a weaker reduction, indicating the need for open efforts in further research. Regarding inference methods, MemFree decoding, which avoids $n$-gram copying from copyright-protected data when determining the next token, successfully reduces literal copying but does not reduce non-literal copying (Figure 2-c). These results highlight an urgent need to study effective mitigation approaches that can alleviate both literal and non-literal reproduction of copyrighted contents while preserving utility. To foster community efforts, we open source our data and code.
@@ 2.1 Legal Framework of Copyright
Literal copying—extensive and verbatim copying without significant alteration—are more likely to be infringing Harper & Row, Publishers, Inc. v. Nation Enterprises (1985). Yet, non-literal copying of an author’s style or the use of similar plots and characters can also infringe Dr. Seuss Enterprises, L.P. v. ComicMix LLC (2020); Paramount Pictures Corp. v. Axanar Productions, Inc. (2017). On the other hand, altering the original work with new expression to change its meaning or message is more likely to be a non-infringing transformative fair use Campbell v. Acuff-Rose Music, Inc. . In addition, facts and ideas are generally not copyrightable, and the allowable scope of copying from a primarily factual work is greater Feist Publications, Inc. v. Rural Telephone Service Co., Inc. (1991). Therefore, it is beneficial for AI systems to memorize and utilize these facts in language model outputs.
@@
However, it is generally accepted that the heart of determination lies in the extent of similarity between the two works Rebikoff (2001); Henderson et al. (2023). In a famous article, legal scholar Zechariah Chaffee identified “the sequence of events and the development of the interplay of the characters” as “the pattern of the work” that is subject to protection Chafee (1945). As one court described it, “the essence of a novel or any other story for that matter, is the plot, plan, arrangement, characters and dialogue therein contained” Bartels (1965). Inspired by these court records, we evaluate non-literal copying by identifying the production of events and characters in the creative writing of language models.
@@ 2.2 Evaluating and Mitigating Copyright Risks in LMs
Most research on LM copyright evaluation has focused on literal copying, i.e., analyzing model outputs for near-exact overlaps with copyrighted snippets Henderson et al. (2023); Meeus et al. (2024); Ippolito et al. (2023); Carlini et al. (2023). However, to bridge the gap to real-world practices, it is necessary to study higher-level semantic similarities. Lee et al. (2023a) evaluate the replication of paraphrases and ideas in language model output within the context of plagiarism rather than copyright violation. Paraphrasing facts generally carries fewer copyright risks. In contrast, our focus is on copyright issues by examining the reproduction of events and characters in creative writing. Additionally, Lu et al. (2024) explored semantic similarities in multi-modal settings, emphasizing the replication of symbols, content, and style in image generation.
@@ 3 CopyBench: Evaluating Reproduction of Copyrighted Text
We introduce CopyBench, a benchmark that provides automatic evaluation of the reproduction of popular copyright-protected fictions as well as the utility of the model. In particular, we evaluate two types of reproduction: literal and non-literal copying. To the best of our knowledge, our work is the first that evaluates non-literal reproduction of copyrighted work in language model generation.
@@ 3.1 Overview
We consider two types of similarity between the LM output and text sources 1. Literal copying occurs when a model’s output contains near-identical portions of the text source. In contrast, non-literal copying occurs when a model’s outputs are similar to the text source at a higher level of abstraction, even if they are not word-for-word identical. This is to evaluate the extent to which a generated story is original when the language model is prompted to write an original story. Although whether the story is original or not is highly context dependent, in this paper, we consider the similarity in events and characters of a story, inspired by piror work in copyright protection of literary work Henderson et al. (2023).
@@ 3.2 Source Text Collection
Our evaluation pipeline can be applied to various sources of copyrighted works. In CopyBench, we focus on fictional books Meeus et al. (2024); Chang et al. (2023); Shi et al. (2023). For literal copying, we randomly sampled snippets from popular copyright-protected fiction. To minimize copyright risks, we choose not to additionally release the actual texts of these copyrighted books. Instead, we created our dataset using existing datasets and included 16 books from BookMIA Shi et al. (2023), which are likely in ChatGPT’s training data as suggested by Chang et al. (2023).
For non-literal copying, we identify 118 fictions in CliffNotes study guide, where each novel is associated with a human-written summary. [^2] We used different sources for literal and non-literal copying because exact snippets are not available in CliffNotes and summaries are not available in BookMIA. Following Chang et al. (2023), we exclude non-fiction books and books whose copyrights have expired (published prior to 1923), only leaving text sources that are copyrighted at the time of writing this paper.
@@ 3.3 Evaluation Tasks and Metrics
We measure non-literal copying in the context of creative writing to evaluate the extent to which a generated story is original with respect to events and characters, following prior court cases. We extract a list of key events and character names from the Cliffnotes summary of the book (see Section A.3 for details). We prompt LMs to generate an original story given the beginning of a story from one of the events in the list.
@@
We extract character names and aliases from the summary because matching characters solely by their full names often leads to many misses. Character Overlap is identified through exact matches of character aliases. If any alias of a character is recalled, the character is considered recalled. To prevent excessive non-literal copying, we exclude characters whose names appear in the prompt. We report the proportion of instances where character overlap exceeds a threshold. For characters, the threshold is set at 3, which is relatively large since the average number of characters in the reference list is 9.0.
@@
We calculate the accuracy, precision, recall, and F1 score for each individual event attribution. Additionally, we compute the Pearson correlation between the number of overlapping events identified by the two methods. The results are presented in . The human evaluation achieves an F1 score of 0.76 and a correlation of 0.70, which are considered high for attribution models.
@@
3.4 Human Analysis of Automatic Event Copying Evaluation
To verify the alignment between automatic event copying evaluation and human judgment, we conducted a human study. We collected outputs from three models: Llama2-70B, Llama2-70B-Chat, and Tulu2-70B. To cover a range of similarity levels between the generated stories and the original works, we selected 10 samples for each automatic event overlap score. If fewer than 10 samples were available for a given score, we included all available samples, resulting in a total of 82 cases. For human annotation, we provided a list of reference events and asked the annotator to determine whether each event was included in the generated story.
We reported the accuracy, precision, recall and F1 score, treating it as a binary classification task in Table 3. Additionally, we reported the Pearson correlation between the event overlap scores (ranging from 0 to 20) obtained through the automatic method and human annotation. Our automatic methods achieved an F1 score of 0.76, which is comparable to the best performance range of 0.6 to 0.85 F1 reported on various attribution datasets Li et al. (2024b). Our event copying evaluation effectively measures the similarity between language model-generated stories and reference fictions based on given events.
Accuracy Precision Recall F1 Correlation
0.89 0.68 0.87 0.76 0.70
@@ 6 Conclusion
No legal conclusions should be drawn from our experiments. Nevertheless, we hope our methods and results provide empirical data to ground discussions in copyright issues for language models.
@@ Acknowledgements
We express our gratitude to Weijia Shi for the fruitful discussions during the early stages of this project. We also thank Rulin Shao, Hamish Ivison, Rui Xin, and Yizhong Wang for their insightful feedback. Finally, we like to thank A. Feder Cooper and Katherine Lee for for helpful comments. PWK is supported by the Singapore National Research Foundation and the National AI Group in the Singapore Ministry of Digital Development and Innovation under the AI Visiting Professorship Programme (award number AIVP-2024-001). This research is also supported in part by Darpa ITM grant ONR N00014-24-1-2207.
@@ A.4 Human Evaluation for Automatic Event Copying Evaluation
The annotators are asked to read the LM-generated story and decide whether each provided reference event is entailed in the story. The instruction is shown in Table 10. We then reported the accuracy, recall and precision of the automatic event detection model and analyzed the correlation between human evaluations and the automatic scores.
@@ C.2 Case Study
2. Review the 20 key events from a well-known copyrighted fictional book.
3. For each key event, determine if it occurs in the generated story by using 1 to indicate the event occurs and 0 if it does not.
Table 10: Instruction for human evaluation for automatic event overlap detection.

### v2

# EMNLP 2024 Main Conference CopyBench: Measuring Literal and Non-Literal Reproduction of Copyright-Protected Text in Language Model Generation

@@ Abstract
Tong Chen1Akari Asai1*Niloofar Mireshghallah1* Sewon Min1James Grimmelmann2,3Yejin Choi1,4 Hannaneh Hajishirzi1,4Luke Zettlemoyer1Pang Wei Koh1,4 1University of Washington   2Cornell University   3Cornell Law School  4Allen Institute for AI
@@ 1 Introduction
To bridge the gap, we introduce CopyBench, a new benchmark and automatic evaluation protocols designed to assess reproduction of copyright-protected text by LMs (Figure 1). We evaluate two categories of copying: literal and non-literal copying. Literal copying assesses the extent to which a model can reproduce copyright-protected content exactly as it appears in the source material. In contrast, non-literal copying evaluates whether a model generates outputs that, despite differing in surface form (e.g., through paraphrasing), exhibit a high degree of overlap in content. To the best of our knowledge, our work is the first that evaluates non-literal reproduction of copyrighted works in language model generation. In order to study the trade-offs between the unintended copying and the desired utilities of LMs, we also quantify two aspects of desired utilities: fact recall, i.e., answering questions about book content, and fluency. Our benchmark therefore allows for the evaluation of levels of copyright work reproduction, overall model utility, and any associated trade-offs. We curate a dataset using a list of popular copyright-protected fictions, sourced from the famous CliffsNotes study guides, [^1] https://www.cliffsnotes.com/ which provide human-written plot summary for each book.
@@
Additionally, our datasets are designed to benchmark methods for potentially reducing copying behavior, broadly categorized into training methods (instruction tuning and chat alignment) and inference methods (e.g., MemFree decoding; Ippolito et al. 2023). We find that different instruction-tuning methods have varying levels of effectiveness. Specifically, Llama2-Chat Touvron et al. (2023) and Llama3-Instruct AI@Meta (2024) significantly reduce copying behavior, though the mechanism remains unclear due to the use of closed-source data. In contrast, the open-source model Tulu2 Ivison et al. (2023), which is based on Llama2 and further trained with fully open-sourced instruction tuning and preference data, shows a weaker reduction, indicating the need for open efforts in further research. Regarding inference methods, MemFree decoding, which avoids $n$-gram copying from copyright-protected data when determining the next token, successfully reduces literal copying but does not reduce non-literal copying (Figure 2-c). These results highlight an urgent need to study effective mitigation approaches that can alleviate both literal and non-literal reproduction of copyrighted contents while preserving utility. To foster community efforts, we open source our data and code.
@@ 2.1 Legal Framework of Copyright
Literal copying—extensive and verbatim copying without significant alteration—is more likely to be infringing Harper & Row, Publishers, Inc. v. Nation Enterprises (1985). Yet, non-literal copying of an author’s style or the use of similar plots and characters can also infringe Dr. Seuss Enterprises, L.P. v. ComicMix LLC (2020); Paramount Pictures Corp. v. Axanar Productions, Inc. (2017). On the other hand, altering the original work with new expression to change its meaning or message is more likely to be a non-infringing transformative fair use Campbell v. Acuff-Rose Music, Inc. . In addition, facts and ideas are generally not copyrightable, and the allowable scope of copying from a primarily factual work is greater Feist Publications, Inc. v. Rural Telephone Service Co., Inc. (1991). Therefore, it is generally beneficial for AI systems to learn and utilize facts in LM outputs.
@@
It is generally accepted that the heart of determination lies in the extent of similarity between the two works Rebikoff (2001); Henderson et al. (2023). In a famous article, legal scholar Zechariah Chaffee identified “the sequence of events and the development of the interplay of the characters” as “the pattern of the work” that is subject to protection Chafee (1945). As one court described it, “the essence of a novel or any other story for that matter, is the plot, plan, arrangement, characters and dialogue therein contained” Bartels (1965). Inspired by these analyses, we evaluate non-literal copying by identifying the production of events and characters in creative-writing outputs of language models.
@@ 2.2 Evaluating and Mitigating Copyright Risks in LMs
Most research on LM copyright evaluation has focused on literal copying, i.e., analyzing model outputs for near-exact overlaps with copyrighted snippets Henderson et al. (2023); Meeus et al. (2024); Ippolito et al. (2023); Carlini et al. (2023). However, to bridge the gap to real-world practices, it is necessary to study higher-level semantic similarities.
Lee et al. (2023a) evaluate the replication of paraphrases and ideas in language model output through the lens of plagiarism rather than copyright infringement. But plagiarism is primarily an ethical academic rule about giving proper attribution for the use of other people’s ideas, while copyright infringement is a legal rule against using an author’s particular expression of ideas Green (2002); Cyphert (2023). Paraphrasing facts from a source might constitute plagiarism but not copyright infringement; copying long passages with proper attribution might be copyright infringement but not plagiarism. The two require different analyses and benchmarks.
Lu et al. (2024) and He et al. (2024) explored semantic similarities in image generation, emphasizing the replication of symbols, content, and style in image generation. Our work, in contrast, focuses on text outputs from LMs.
Concurrently with our work, Wei et al. (2024) and Liu et al. (2024) also study copyright risks in LM generation. However, their focus is primarily on benchmarking literal copying and its mitigation methods, whereas our work explores non-literal copying as well.
@@
2.3 Memorization and Extraction in LMs
The root cause of copying behaviour is the memorization of training data in LMs, and then its regurgitation in a provided context i.e. in response to given prompts Biderman et al. (2024); Carlini et al. (2021); Nasr et al. (2023). Recent work studying memorization has profiled patterns, such the impact that the number of gradient updates, batch size, the length of a sample, and the size of the model have on its memorization Huang et al. (2024); Prashanth et al. (2024); Duan et al. (2024); Mireshghallah et al. (2022). These works however, only study memorization in the original context that the data point appeared in, in the training corpus. They also focus on base models and not instruction-tuned ones. Closely related to ours are Schwarzschild et al. (2024); Kassem et al. (2024), which take an adversarial approach, by optimizing prompts for the sake of maximizing extraction. These works also look into instruction tuned models and reveal higher levels of memorization through prompt optmization. Although our work is related to memorization and extraction literature, the goals and assumptions are slightly different: we are targeting copyrighted material, and assessing risk with respect to the surrounding legal framework. Extraction attacks are primarily focused on retrieving training data, regardless of its legality.
@@ 3 CopyBench: Evaluating Reproduction of Copyrighted Text
We introduce CopyBench, a benchmark that provides automatic evaluation of the reproduction of popular copyright-protected fictions as well as the utility of the model. In particular, we evaluate two types of reproduction: literal and non-literal copying. To the best of our knowledge, our work is the first that evaluates non-literal reproduction of copyrighted work in LMs.
@@ 3.1 Overview
We consider two types of similarity between the LM output and text sources Figure 1. Literal copying occurs when a model’s output contains near-identical portions of the text source. In contrast, non-literal copying occurs when a model’s outputs are similar to the text source at a higher level of abstraction, even if they are not word-for-word identical. This is to evaluate the extent to which a generated story is original when the language model is prompted to write an original story. Although whether the story is original or not is highly context dependent, in this paper, we consider the similarity in events and characters of a story, inspired by prior work in copyright protection of literary work Henderson et al. (2023).
@@ 3.2 Source Text Collection
Our evaluation pipeline can be applied to various sources of copyrighted works. In CopyBench, we focus on fiction books Meeus et al. (2024); Chang et al. (2023); Shi et al. (2023). For literal copying, we randomly sampled snippets from popular copyright-protected fiction. To minimize copyright risks, we choose not to additionally release the actual texts of these copyrighted books. Instead, we created our dataset using existing datasets and included 16 books from BookMIA Shi et al. (2023), which are likely in ChatGPT’s training data as suggested by Chang et al. (2023).
For non-literal copying, we identify 118 fictions in CliffsNotes study guide, where each novel is associated with a human-written summary. [^2] We used different sources for literal and non-literal copying because exact snippets are not available in CliffsNotes and summaries are not available in BookMIA. Following Chang et al. (2023), we exclude non-fiction books and books whose copyrights have expired (published prior to 1923), only leaving text sources that are copyrighted at the time of writing this paper.
@@ 3.3 Evaluation Tasks and Metrics
We measure non-literal copying in the context of creative writing to evaluate the extent to which a generated story is original with respect to events and characters, following prior court cases. We extract a list of key events and character names from the CliffsNotes summary of the book (see Section A.3 for details). We prompt LMs to generate an original story given the beginning of a story from one of the events in the list.
@@
We extract character’s names (including first name, last name and aliases) from the summary because matching characters solely by their full names often leads to many misses. Character Overlap is identified through exact matches of character names. If any name of a character is recalled, the character is considered recalled. To prevent excessive non-literal copying, we exclude characters whose names appear in the prompt. We report the proportion of instances where character overlap exceeds a threshold. For characters, the threshold is set at 3. For comparison, the average number of characters in the reference list across books in our dataset is 9.0.
@@
3.4 Human Verification of Automatic Evaluation
To verify that the automatic metrics align with human understanding of event and character reproduction, we conduct a human annotation on a subset of test cases. We provide the annotators with the generated story along with a list of events and characters, asking them to identify which of these are present in the story. Each event or character is treated as a binary classification problem (included or not), allowing us to compute accuracy, precision, recall, and F1 score, as shown in Table 3. The F1 score for event overlap is 0.76, comparable to state-of-the-art methods on other attribution datasets Li et al. (2024b), indicating good performance. The F1 score for character overlap is 0.96, approaching the perfect score of 1.0. These human annotation results demonstrate high accuracy for the automatic metric.
Accuracy Precition Recall F1
Event Overlap 0.89 0.68 0.87 0.76
Character Overlap 0.96 0.93 0.95 0.94
@@ 6 Conclusion
No legal conclusions should be drawn from our experiments. Our methods shed light on ways to understand overall behavioral properties of LM systems, and are not suitable for making any specific assessments of infringement or noninfringement. Nevertheless, we hope our methods and results provide empirical data to ground discussions on copyright issues for language models.
@@ Acknowledgements
We express our gratitude to Weijia Shi for the fruitful discussions during the early stages of this project. We also thank Rulin Shao, Hamish Ivison, Rui Xin, and Yizhong Wang for their insightful feedback. Finally, we like to thank A. Feder Cooper and Katherine Lee for for helpful comments. PWK is supported by the Singapore National Research Foundation and the National AI Group in the Singapore Ministry of Digital Development and Innovation under the AI Visiting Professorship Programme (award number AIVP-2024-001). YC is supported in part by Darpa ITM grant ONR N00014-24-1-2207.
@@ A.4 Human Evaluation for Automatic Event Copying Evaluation
The annotators are asked to read the LM-generated story and decide whether each provided reference event and character is entailed in the story. The instruction is shown in Table 11. We then reported the accuracy, recall and precision of the automatic event detection model.
@@ C.2 Case Study
Rating Generated Text Reason
1 […] O’Brien invites him to his house. Winston and Julia visit O’Brien. O’Brien gives them a copy of Goldstein’s book. Winston reads the book. Winston and Julia meet in the room above Mr Charrington’s shop. […] The story lacks coherence and is hard to comprehend. It seems to be a list of events, rather than a narrative.
5 […] He is given a sword named Sting, which glows blue whenever Orcs are nearby. Bilbo is initially reluctant to accept the weapon, but he eventually takes it. The group travels through the Misty Mountains, where they encounter three trolls. […] The story is written in a clear and organized manner.
Table 10: Fluency evaluation of the generated text.
@@
2. Review the events and character aliases from a well-known copyrighted fictional book.
3. For each event or character, determine if it occurs in the generated story by using 1 to indicate the it occurs and 0 if it does not.
Table 11: Instruction for human evaluation for automatic overlap evaluation.